OpenStudy (anonymous):

If Larry has 250 marbles, all red , white, and blue, in the ratio of 1:3:6. How many blue marbles does he have???

OpenStudy (amistre64):

6+3+1 = 10*25 = 250 6*25

OpenStudy (anonymous):

1x+3x+6x=250 10x=250 x=250/10 x=25 we took x as common factor in the ratio 1x=1*25=25 3x=75 6x=150

OpenStudy (anonymous):

thx